bill533:

Lars rides a chairlift to the top ofa mountain. The chairlift rises at a constant angle of 37°. If the length of the chairlift ride is 1,392 m, what is the elevation gain from the base of the chairlift to the top? Draw a right triangle to model this problem and use the Gizmo to find sin 37°. Show your work. Elevation gain:

bill533:

@jhonyy9

bill533:

i believe the answer is 3/5

bill533:

1 attachment
jhonyy9:

drawn this right triangle please

jhonyy9:



bill533:

1 attachment
jhonyy9:

sin 37 = BC / 1392 yes ?

jhonyy9:

sin 37 = opposite side divide hypothenuse

bill533:

what about the question 5

jhonyy9:

what is q5 ?

bill533:

it's on the picture i posted

jhonyy9:

ok. so this is easy just use the sin theorem

jhonyy9:

sin 12 = oppo / hyp

jhonyy9:

you dont calcule it ? why ?

bill533:

In ΔDEF, <F is a right angle. Suppose m<D = 12° and EF = 1.3. Follow the steps below to find the length of the hypotenuse, DE. A. To solve this problem, first draw a sketch of ΔDEF in the space to the right. B. You know that sin D = OPP/HYP. In the space to the right, substitute the known values into this equation. Use x for the unknown, DE. C. Use algebra to solve for x. Show your work to the right. Then, use the Gizmo to find the sine value you need. Finally, find DE with a calculator. What is the length of DE?

jhonyy9:

sin 12 = 1,3 / DE DE = ?

bill533:

9.2

jhonyy9:

DE = 1,3 / (sin 12°) = ?

jhonyy9:

sin 12° = 0,207 so equal 0,21 DE = 1,3/0,21 = 6,19 hope helped

bill533:

0.20

jhonyy9:

why 0,20 ? the 3rd term is greater than 5 so we can considering 0,21

bill533:

it was for sin 12

jhonyy9:

yes exactly

bill533:

sorry i took long to answer, i had many tabs

bill533:

ok i got that, back to 6 what was the elevation gain

bill533:

i didn't get that one, it's the top question